# RT-DATR:Real-time Unsupervised Domain Adaptive Detection Transformer with Adversarial Feature Learning

### v1

@@ 3.4. Instance Feature Alignment and Consistency Loss
(5) $q_{i+1}=Linear(MultiHeadAttn(q_{i},k_{i},v_{i}))$
(6) $\mathcal{L}^{D}_{adv}=-dlog(D_{d}(q_{i})+(1-d)log(D_{d}(q_{i}))$
Herein, $q$ denotes the domain query, $i$ represents the layer index of the decoder ($i=0,1,...,N$), while $k$ and $v$ are both provided by the object query. $D_{d}$ represents the discriminator within the decoder.
@@
(7) $\mathcal{L}_{cons}=\frac{1}{N}\sum_{i=1}^{N}JSD(\hat{y}_{cls},y^{i}_{cls})$
@@ 3.5. Total Loss
(8) $\mathcal{L}_{total}=\mathcal{L}_{det}+\lambda_{1}\mathcal{L}^{B}_{adv}+\lambda_{2}\mathcal{L}^{E}_{adv}+\lambda_{3}\mathcal{L}^{D}_{adv}+\lambda_{4}\mathcal{L}_{cons}$
@@
4. Experiments
We evaluated our approach on multiple scene datasets, including weather adaptation (Cityscapes (Cordts et al., 2016) to Foggy Cityscapes (Sakaridis et al., 2018)), scene adaptation (Cityscapes to BDD100K (Yu et al., 2020)), artistic-to-real adaptation (Sim10K (Johnson-Roberson et al., 2016) to Cityscapes) and cross-camera adaptation (KITTI (Geiger et al., 2012) to Cityscapes). Ablation experiments were also conducted on the modules proposed in this paper. All results indicate that the approach presented in this paper is highly adaptable to various domain adaptation tasks across different scene datasets, significantly improving the algorithm’s generalization performance.
Table 2. Experimental results of different methods for scene adaptation, Cityscapes → BDD100k daytime subset. R34 and R50 represent ResNet-34 and ResNet-50, respectively. V-16 represents VGG-16. Def DETR represents Deformable DETR.
@@ 4. Experiments
SFA (Wang et al., 2021) R50 40.2 27.6 57.5 19.1 23.4 15.4 19.2 28.9
AQT (Huang et al., 2022) R50 38.2 33.0 58.4 17.3 18.4 16.9 23.5 29.4
MTTrans(Yu et al., 2022) R50 44.1 30.1 61.5 25.1 26.9 17.7 23.0 32.6
MRT (Zhao et al., 2023) R50 48.4 30.9 63.7 24.7 25.5 20.2 22.6 33.7
MTM (Weng and Yuan, 2024) R50 53.7 35.1 68.8 23.0 28,8 23.8 28.0 37.3
Source Only (RT-DETR) R34 47.5 37.4 69.7 27.0 28.0 24.9 32.3 33.3
Source Only (RT-DETR) R50 51.6 40.3 71.5 31.6 35.3 29.0 34.3 36.7
RT-DATR(Ours) R34 54.0 44.3 73.0 33.6 35.6 30.4 34.9 38.2
RT-DATR(Ours) R50 57.0 45.9 74.5 36.4 39.7 33.2 37.9 40.7
4.1. Datasets
4.1.1. Cityscapes to FoggyCityscapes
Cityscapes is a semantic understanding dataset. It contains 2,975 training images and 500 validation images. FoggyCityscapes is a dataset generated by adding varying degrees of fog to the original Cityscapes images. The fog conditions are categorized into three levels: 0.005, 0.01, and 0.02, representing light, medium, and heavy fog, respectively. We selected the most challenging level, 0.02, to evaluate our experimental results on 8 common categories.
4.1.2. Cityscapes to BDD100K
BDD100K is a large-scale autonomous driving dataset that includes 70,000 training images and 10,000 validation images. We selected a subset that contains only daytime images to construct the adaptation benchmark. Therefore, we used 36,728 images for training and 5,258 images for validation. We chose Cityscapes as the source domain data and BDD100K as the target domain data, and evaluated 7 common categories.
4.1.3. Sim10K to Cityscapes
Sim10K is a dataset generated by a game engine, which contains 10,000 images and corresponding detection boxes. In the experiment, we selected Sim10K as the source domain data and Cityscapes as the target domain data. Only the “car” category was included in the evaluation.
Figure 3. Visualization of results on BDD100K using different approaches.
4.1.4. KITTI to Cityscapes
KITTI, another street scene dataset, differs from Cityscapes in that it originates from different cities and cameras. For our experiments, we utilized 7481 images as the training set, focusing solely on the car category that is common to both KITTI and Cityscapes.
4.2. Implemention Details
We have selected RT-DETR as our base detector, utilizing ResNet-34 and ResNet-50 as backbone networks. The learning rate is set at 2e-4, and the batch size is 2. Training is performed on 8 A100 GPUs. For other settings, we follow the RT-DETR configuration, using the AdamW optimizer, 72 training epochs, and specifying an inference size of 640x640. Regarding loss weight coefficients, all are set to 1, except for $\lambda_{1}$, which is assigned a value of 1.5. For evaluation metrics, we report the Average Precision (AP) for each object category and the mean Average Precision (mAP) across all categories, with an IoU threshold of 0.5. Additionally, to ensure reliable initialization weights for the target domain, we pretrain the model on the source domain dataset.
@@ 4.2. Implemention Details
Table 3. Experimental results of the scenario synthetic scene to real scene: SIM10k → Cityscapes. R34 and R50 represent ResNet-34 and ResNet-50, respectively. Def DETR represents Deformable DETR.
@@
DA-Faster (Chen et al., 2018) R50 41.9
GPA (Xu et al., 2020a) R50 47.6
KTNet (Tian et al., 2021) R50 50.7
PT (Tian et al., 2021) R50 55.1
@@
SFA (Wang et al., 2021) R50 52.6
AQT (Huang et al., 2022) R50 53.4
DA-DETR (Zhang et al., 2023) R50 54.7
MRT (Zhao et al., 2023) R50 62.0
MTM (Weng and Yuan, 2024) R50 58.1
@@
4.3. Comparison with Other Methods
In this section, we compare the detection performance of RT-DATR with other current state-of-the-art methods (based on Faster R-CNN, FCOS, and DETR-style detectors) across four domain adaptation scenarios. It is important to note that due to the architectural differences between one-stage and two-stage detectors, as well as the varying module designs in DETR-style detectors, directly transferring other methods to our baseline for comparison may not be entirely fair. Both AQT and ConfMix conducted ablation studies and discussions on the impact of different detector components on model performance. Therefore, we aim to further push the absolute performance of the model while emphasizing its real-time capability rather than opting for a lower baseline. This move represents progress for both academic and industrial applications. Additional analysis is provided in Appendix A.
4.3.1. Weather Adaptation
We first studied the adaptation from normal weather to foggy conditions based on the task Cityscapes→ Foggy Cityscapes. As shown in Table 1, RT-DATR outperforms both the baseline and other state-of-the-art methods. For instance, with a ResNet34 backbone, RT-DATR achieves an mAP of 50.4%, which is comparable to the metrics of models using a student-teacher framework, typically considered to yield better results. When using a ResNet50 backbone, RT-DATR attains an mAP of 52.7%, surpassing the existing SOTA method, CAT, by 0.2%. Furthermore, in specific categories such as “ride”, “car”, “truck”, RT-DATR demonstrates higher AP metrics compared to other methods.
4.3.2. Scene adaptation
Table 2 presents the results of the cross-scene adaptation experiment for the Cityscapes→BDD100K task. Following (Xu et al., 2020b), we report results for 7 common categories. It is evident that RT-DATR outperforms the most recent domain adaptive object detection methods, enhancing the mAP from 36.7% achieved by the baseline model to 40.7%. Notably, RT-DATR outperforms both the baseline and other state-of-the-art methods across all 7 categories.
4.3.3. Artistic to Real Adaptation
Training a target detector from synthetic images and generalizing it to the real world can significantly enhance the diversity and utility of data sources. Table 3 presents the artistic-to-real adaptation experiment for the task Sim10K→Cityscapes. RT-DATR achieves an AP of 67.2%, which is 5.2% higher than the current state-of-the-art method, MRT. Furthermore, compared to CNN-based detectors, RT-DATR exhibits significant advantages in performance while also possessing real-time capabilities.
Table 4. Experimental results of the cross-camera adaptation: KITTI → Cityscapes. R34 and R50 represent ResNet-34 and ResNet-50, respectively. V16 represents VGG-16. Def DETR represents Deformable DETR.
@@ 4.3.3. Artistic to Real Adaptation
4.3.4. Cross-camera Adaptation
Table 4 presents the cross-camera adaptation experiments for the task of KITTI→Cityscapes. It can be observed that RT-DATR achieves an AP of 50.3% when utilizing a ResNet34 backbone, outperforming DA-DETR by +1.4% and TIA by +6.3%. These experimental results further demonstrate that the proposed method can generalize well to different domain adaptation tasks.
Table 5. Quantitative ablation study of RT-DATR on Cityscapes to FoggyCityscapes. R34 represents ResNet-34.
@@ 4.3.4. Cross-camera Adaptation
4.3.5. Real-time Analysis
It is worth noting that our endeavor does not lie in creating a novel real-time object detector. Rather, we strive to enhance the cross-domain detection capabilities of an existing transformer-based real-time detector, particularly emphasizing practical applications—a realm that remains largely uncharted in current research. The domain adaptation modules proposed in RT-DATR are training-only, ensuring that its inference latency remains consistent with that of RT-DETR. We conducted latency testing using TensorRT FP16 on a T4 GPU, with RT-DATR-R34 achieving a result of 6.3 ms. We believe that this real-time domain adaptive detection transformer has significant potential for widespread application.
4.4. Ablation Study
@@
4.5. Visualization
We utilize t-SNE (Van der Maaten and Hinton, 2008) to visualize the features extracted from both the backbone and encoder on Cityscapes to Foggy Cityscapes. As illustrated in Figure 4, compared to the relatively dispersed feature distribution between the source and target domains in RT-DETR, the features in RT-DATR are more indistinguishable. Furthermore, the encoder features in RT-DATR exhibit a higher degree of clustering compared to the backbone features, demonstrating that our method is capable of learning domain-invariant feature representations. Furthermore, we compared the detection performance of the model, as shown in Figure 3. RT-DETR and MRT exhibit similar performance, while RT-DATR outperforms MRT and RT-DETR, achieving a higher precision and recall rate.
Figure 4. Feature visualization of the two domains on Cityscapes to Foggy Cityscapes by t-SNE. The blue and red points denotes source and target features respectively.
5. Conclusion
In this work, we propose a real-time domain-adaptive detection transformer, RT-DATR. We achieve local object-level and scene semantic feature alignment through adversarial feature learning, further mitigating the domain gap with domain queries, and enhance the model’s robustness through consistency loss. Experimental results on multiple cross-domain benchmarks demonstrate the effectiveness of our method. Furthermore, considering the real-time capabilities and state-of-the-art performance of RT-DATR, we believe this can offer new insights and support to both the academic and industrial communities. In the future, we aim to explore various domain adaptation strategies based on real-time detection transformers, with the objective of further enhancing cross-domain real-time detection performance.

### v2

@@ 3.4 Instance Feature Alignment and Consistency Loss
$q_{i+1}=Linear(MultiHeadAttn(q_{i},k_{i},v_{i}))$ (5)
$\mathcal{L}^{D}_{adv}=-dlog(D_{d}(q_{i})+(1-d)log(1-D_{d}(q_{i}))$ (6)
Herein, $q$ denotes the domain query, $i$ represents the layer index of the decoder (i=0,1,,Ni=0,1,...,N), while $k$ and $v$ are both provided by the object query. $D_{d}$ represents the discriminator within the decoder.
@@
$\mathcal{L}_{cons}=\frac{1}{N}\sum_{i=1}^{N}JSD(\hat{y}_{cls},y^{i}_{cls})$ (7)
@@ 3.5 Total Loss
$\mathcal{L}_{total}=\mathcal{L}_{det}+\lambda_{1}\mathcal{L}^{B}_{adv}+\lambda_{2}\mathcal{L}^{E}_{adv}+\lambda_{3}\mathcal{L}^{D}_{adv}+\lambda_{4}\mathcal{L}_{cons}$ (8)
@@
SFA [31] R50 40.2 27.6 57.5 19.1 23.4 15.4 19.2 28.9
AQT [15] R50 38.2 33.0 58.4 17.3 18.4 16.9 23.5 29.4
MTTrans[37] R50 44.1 30.1 61.5 25.1 26.9 17.7 23.0 32.6
MRT [42] R50 48.4 30.9 63.7 24.7 25.5 20.2 22.6 33.7
BiADT [12] R50 42.0 34.5 59.9 17.2 19.2 17.8 24.4 32.7
MTM [32] R50 53.7 35.1 68.8 23.0 28.8 23.8 28.0 37.3
Source Only (RT-DETR) R34 47.5 37.4 69.7 27.0 28.0 24.9 32.3 38.1
Source Only (RT-DETR) R50 51.6 40.3 71.5 31.6 35.3 29.0 34.3 41.9
RT-DATR(Ours) R34 54.0 44.3 73.0 33.6 35.6 30.4 34.9 43.7
RT-DATR(Ours) R50 57.0 45.9 74.5 36.4 39.7 33.2 37.9 46.5
@@
Table 2: Experimental results of different methods for scene adaptation, Cityscapes → BDD100k daytime subset. R34 and R50 represent ResNet-34 and ResNet-50, respectively. V-16 represents VGG-16. Def DETR represents Deformable DETR.
4 Experiments
We evaluated our approach on multiple cross-domain datasets, including weather adaptation (Cityscapes [5] to Foggy Cityscapes [25]), scene adaptation (Cityscapes to BDD100K [36]), artistic-to-real adaptation (Sim10K [17] to Cityscapes) and cross-camera adaptation (KITTI [10] to Cityscapes). Ablation experiments were also conducted on the modules proposed in this paper. All results indicate that the approach presented in this paper is highly adaptable to various domain adaptation tasks across different scene datasets, significantly improving the algorithm’s generalization performance.
4.1 Datasets
Cityscapes to FoggyCityscapes. Cityscapes is a semantic understanding dataset. It contains 2,975 training images and 500 validation images. FoggyCityscapes is a dataset generated by adding varying degrees of fog to the original Cityscapes images. The fog conditions are categorized into three levels: 0.005, 0.01, and 0.02, representing light, medium, and heavy fog, respectively. We selected the most challenging level, 0.02, to evaluate our experimental results on 8 common categories.
Cityscapes to BDD100K. BDD100K is a large-scale autonomous driving dataset that includes 70,000 training images and 10,000 validation images. We selected a subset that contains only daytime images to construct the adaptation benchmark. Therefore, we used 36,728 images for training and 5,258 images for validation. We chose Cityscapes as the source domain data and BDD100K as the target domain data, and evaluated 7 common categories.
Sim10K to Cityscapes. Sim10K is a dataset generated by a game engine, which contains 10,000 images and corresponding detection boxes. In the experiment, we selected Sim10K as the source domain data and Cityscapes as the target domain data. Only the “car” category was included in the evaluation.
KITTI to Cityscapes. KITTI, another street scene dataset, differs from Cityscapes in that it originates from different cities and cameras. For our experiments, we utilized 7481 images as the training set, focusing solely on the car category that is common to both KITTI and Cityscapes.
4.2 Implemention Details
We selected RT-DETR as our base detector, utilizing ResNet-34 and ResNet-50 as backbone networks. The learning rate is set at 2e-4, and the batch size is 2. Training is performed on 8 A100 GPUs. For other settings, we follow the RT-DETR configuration, using the AdamW optimizer, 72 training epochs, and specifying an inference size of 640x640. Regarding loss weight coefficients, all are set to 1, except for $\lambda_{1}$, which is assigned a value of 1.5. For evaluation metrics, we report the Average Precision (AP) for each object category and the mean Average Precision (mAP) across all categories, with an IoU threshold of 0.5. Additionally, to ensure reliable initialization weights for the target domain, we pretrain the model on the source domain dataset.
4.3 Comparison with Other Methods
In this section, we compare the detection performance of RT-DATR with other current state-of-the-art methods (based on Faster R-CNN, FCOS, and DETR-style detectors) across four domain adaptation scenarios.
Method Faster RCNN Def-DETR RT-DETR mAP
Backbone VGG-16 ResNet-50 ResNet-34
PLA 37.3
PLA 41.8
LOFA(our) 45.6
ILA 36.8
ILA 38.1
IFA(our) 40.9
Table 3: Transfer the existing domain alignment method to RT-DETR and compare it with ours on Cityscapes→Foggy Cityscapes. PLA represnet pixel-level alignment, ILA represent instance-level alignment.
@@ 4.3 Comparison with Other Methods
DA-Faster [4] R50 41.9
GPA [34] R50 47.6
KTNet [26] R50 50.7
PT [26] R50 55.1
@@
SFA [31] R50 52.6
AQT [15] R50 53.4
DA-DETR [39] R50 54.7
MRT [42] R50 62.0
MTM [32] R50 58.1
@@
Table 4: Experimental results of the scenario synthetic scene to real scene: SIM10k → Cityscapes.
Motivation Explanation. It is important to note that due to the architectural differences between one-stage and two-stage detectors, as well as the varying module designs in DETR-style detectors, directly transferring other methods to our baseline for comparison may not be entirely fair. Both AQT and ConfMix conducted ablation studies and discussions on the impact of different detector components on model performance. To further validate this limitation, we transferred the pixel-level alignment (PLA) method from  [43] and the instance-level alignment (ILA) method from AQT to RT-DETR and compared it with our proposed module. As shown in Table 3, our proposed local object-level feature alignment (LOFA) module and instance feature alignment (IFA) module significantly outperform these coarse-grained baseline methods, which confirms the necessity and effectiveness of our work.
Weather Adaptation. We first studied the adaptation from normal weather to foggy conditions based on the task Cityscapes→ Foggy Cityscapes. As shown in Table 1, RT-DATR outperforms both the baseline and other state-of-the-art methods. For instance, with a ResNet34 backbone, RT-DATR achieves an mAP of 50.4%, which is comparable to the metrics of models using a student-teacher framework, typically considered to yield better results. When using a ResNet50 backbone, RT-DATR attains an mAP of 52.7%, surpassing the existing SOTA method, CAT, by 0.2%.
Scene adaptation. Table 2 presents the results of the cross-scene adaptation experiment for the Cityscapes to BDD100K task. Following [33], we report results for 7 common categories. It is evident that RT-DATR outperforms the most recent domain adaptive object detection methods, enhancing the mAP from 41.9% achieved by the baseline model to 46.5%.
Figure 3: Comparison of visualization results using different methods across various cross-domain detection datasets: the BDD100K dataset at the top, the Cityscapes dataset in the middle, and the Cityscapes Foggy dataset at the bottom.
Figure 4: Feature visualization of the two domains on Cityscapes to Foggy Cityscapes by t-SNE. The blue and red points denotes source and target features respectively.
Artistic to Real Adaptation. Training a target detector from synthetic images and generalizing it to the real world can significantly enhance the diversity and utility of data sources. Table 4 presents the artistic-to-real adaptation experiment for the task Sim10K→Cityscapes. RT-DATR achieves an AP of 67.2%, which is 5.2% higher than the current state-of-the-art method, MRT. Furthermore, compared to CNN-based detectors, RT-DATR exhibits significant advantages in performance while also possessing real-time capabilities.
@@
Table 5: Experimental results of the cross-camera adaptation: KITTI → Cityscapes.
Cross-camera Adaptation. Table 5 presents the cross-camera adaptation experiments for the task of KITTI→Cityscapes. It can be observed that RT-DATR achieves an AP of 50.3% when utilizing a ResNet34 backbone, outperforming DA-DETR by +1.4% and TIA by +6.3%. These experimental results further demonstrate that the proposed method can generalize well to different domain adaptation tasks.
Real-time Analysis. It is worth noting that our endeavor does not lie in creating a novel real-time object detector. Rather, we strive to enhance the cross-domain detection capabilities of an existing transformer-based real-time detector, particularly emphasizing practical applications—a realm that remains largely uncharted in current research. The domain adaptation modules proposed in RT-DATR are training-only, ensuring that its inference latency remains consistent with that of RT-DETR. We conducted latency testing using TensorRT FP16 on a T4 GPU, with RT-DATR-R34 achieving a result of 6.3 ms (158 FPS). Not only is it faster, but its performance also surpasses that of the known CNN-based real-time domain adaptive detector, ConfMix. Moreover, we have achieved better results than other non-real-time cross-domain detectors under real-time conditions, and we believe that this real-time domain adaptive detection transformer has significant potential for widespread application.
@@
Table 6: Ablation study of RT-DATR on Cityscapes → Foggy Cityscapes.
4.4 Ablation Study
@@
Hyperparameter Study. We conducted comparative experiments by setting different weight coefficients and discovered that the best performance was achieved when balancing the detection loss and adversarial loss, which is consistent with our prior knowledge.
4.5 Visualization
We use t-SNE [29] to visualize features from the backbone and encoder for the Cityscapes to Foggy Cityscapes adaptation. As illustrated in Figure 4, RT-DATR’s features are less distinguishable between source and target domains compared to RT-DETR’s more dispersed distribution. Furthermore, the encoder features in RT-DATR exhibit a higher degree of clustering compared to the backbone features, demonstrating that our method is capable of learning domain-invariant feature representations. Additionally, we compared the detection performance of the model, as shown in Figure 3. RT-DETR and MRT exhibit similar performance, while RT-DATR outperforms MRT and RT-DETR, achieving a higher precision and recall rate.
5 Conclusion
In this work, we propose a real-time domain-adaptive detection transformer, RT-DATR. We achieve local object-level and scene semantic feature alignment through adversarial feature learning, further mitigating the domain gap with domain queries, and enhance the model’s robustness through consistency loss. Experimental results on multiple cross-domain benchmarks demonstrate the effectiveness of our method. Furthermore, considering the real-time capabilities and state-of-the-art performance of RT-DATR, we believe this can offer new insights and support to both the academic and industrial communities.